OpenStudy (anonymous):

Solve for x. 5x - 11 = 7x - 3

OpenStudy (anonymous):

First you have to isolate x. Do you know how?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Get x by its self by adding from both sides.

OpenStudy (anonymous):

Instagram, that's not really true...

OpenStudy (anonymous):

5x - 11 = 7x - 3 Isolate x by subtracting 7x from both sides. 5x - 7x - 11 = 7x - 3 - 7x

OpenStudy (anonymous):

-2x - 11 = -3 To further isolate x, add 11 to both sides.

OpenStudy (anonymous):

-2x - 11 + 11 = -3 + 11

OpenStudy (anonymous):

-2x = 8

OpenStudy (anonymous):

Now you divide both sides by -2 to get: x = -4

OpenStudy (anonymous):

x=-4 And you can add from both sides because it's subtracting @jcpd910

OpenStudy (anonymous):

"add from both sides" is not very helpful, it's confusing. o.o

OpenStudy (anonymous):

The first step is subtraction.

OpenStudy (anonymous):

But I wasn't wrong :(((

OpenStudy (anonymous):

Haha Thanks from you all

OpenStudy (anonymous):

Whoever gives me a medal gets one

OpenStudy (anonymous):

That's not how medals work.

OpenStudy (anonymous):

That's how you alienate people who helped you, because I obviously helped you a lot more than he did.

OpenStudy (anonymous):

True but you get them all of the time. I need more and i never get them. Thats why i said that.